# Diffusion Models for Time Series Forecasting: A Survey

### v1

@@ 4.2 Diffusion-centric Condition
Diffusion-centric approaches modify the forward process and the reverse process of diffusion models. These approaches follow two primary directions. One research direction integrates time series priors into the forward process, enabling a noise addition procedure specific to time-series data. This prior information also guides learning within the reverse process. The other direction replaces the standard noise addition diffusion with alternative processes, such as a deterministic interpolation process. Such modifications reduce uncertainty inherent in TSF tasks. Both directions represent adaptations of the core diffusion framework to enhance performance for TSF.
Real-world time series inherently possess time-varying uncertainties due to external physical factors. Standard diffusion models enforce a consistent stationary distribution, namely Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$, at the diffusion endpoint for all time series. This motivates recent approaches to incorporate prior knowledge from historical time series within the forward diffusion process. For example, TMDM [61] integrates a Transformer to extract historical features $\hat{\mathbf{y}}_{1:H}$ from $\mathbf{x}_{-L+1:0}$, utilizing this as a conditional prior in both diffusion processes. The modified forward process interpolates between the future time series $\mathbf{x}_{1:H}^{0}$ and the conditional prior $\hat{\mathbf{y}}_{1:H}$ as
@@
In TMDM, the endpoint distribution at diffusion step $T$ is $\mathcal{N}(\hat{\boldsymbol{y}}_{1:H},\mathbf{I})$. The reverse process employs the prior $\hat{\mathbf{y}}_{1:H}$ in the denoising function $\epsilon_{\theta}(\cdot)$ and calculate the $\mathbf{X}_{1:H}^{t-1}$:
$\mathbf{X}_{1:H}^{t-1}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{1:H}^{t}-(1-\sqrt{\bar{\alpha}_{t}})\hat{\mathbf{y}}_{1:H}-{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{1:H}^{t},\hat{\mathbf{y}}_{1:H},t)\right)$ (23)
@@
where $\gamma_{0}$, $\gamma_{1}$, $\gamma_{2}$ and $\tilde{\boldsymbol{\beta}}_{t}$ are parameters that depend only on $\beta_{t}$. This joint conditioning strategy leverages historical dependencies to refine uncertainty estimation in prediction. Building on TMDM, NsDiff [62] further incorporates a historical-dependent variance to model non-stationary uncertainty. The forward process transitions to an endpoint distribution:
@@
$\mathbf{x}_{1:H}^{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{1:H}^{0}+(1-\sqrt{\bar{\alpha}_{t}})\mathbf{f}(\mathbf{x}_{-L+1:0})+\sqrt{(\bar{\beta}_{t}-\tilde{\beta}_{t})\mathbf{g}(\mathbf{x}_{-L+1:0})+\tilde{\beta}_{t}\sigma_{\mathbf{x}_{1:H}}}\epsilon$ (26)
where $\tilde{\beta}_{t}$ is a parameter that depend only on $\alpha_{t}$ and $\sigma_{\mathbf{x}_{1:H}}$ is the actual variance. The reverse process conditions denoising on both $\mathbf{f}(\mathbf{x}_{-L+1:0})$ and $\mathbf{g}(\mathbf{x}_{-L+1:0})$ to refine uncertainty estimation. This approach adapts the diffusion trajectory to temporal uncertainty shifts, improving distributional flexibility. Besides modeling uncertainty, another study integrates recurrent fusion of historical priors within the forward diffusion process to enhance the utilization of recent temporal information. REDI [68] implements a recurrent diffusion mechanism where each forward step combines the current noisy future time series value $\mathbf{x}_{i}^{t}$ at time step $i$ with the immediate historical value $\mathbf{x}_{i-1}^{t}$ before noise addition:
@@
This recurrent formulation strengthens focus on proximate history to mitigate distribution drift and ensures strict temporal causality by preventing information leakage from future states during generation.
Beyond the aforementioned approaches, another research direction focuses on developing alternative formulations for the forward diffusion process. This alternative formulation specifically captures the inherent characteristics of time series data. For instance, DYffusion [63] leverages the temporal dynamics in the data, directly coupling it with the diffusion steps. Specifically, this approach designs a stochastic temporal interpolator $I(\cdot)$ for the forward process to generate intermediate states between the initial condition $\mathbf{x}_{0}$ and the target future point $\mathbf{x}_{H}$, formalized as minimizing $\mathbb{E}\left\|I_{\phi}(\mathbf{x}_{0},\mathbf{x}_{H},i)-\mathbf{x}_{i}\right\|^{2}$ where $i$ is an intermediate timestep. The reverse process employs a forecaster $F(\cdot)$ to predict $\mathbf{x}_{H}$ from interpolated states, defined as minimizing $\mathbb{E}\left\|F(\mathbf{x}_{i},i)-\mathbf{x}_{H}\right\|^{2}$, which enhances computational efficiency and enables flexible continuous-time sampling trajectories. S2DBM [69] proposes a Brownian Bridge diffusion process to replace the traditional Gaussian diffusion framework. This approach constructs a stochastic bridge between a historical prior $\mathbf{h}=F_{p}(\mathbf{x}_{-L+1:0})$ and the target future $\mathbf{x}_{1:H}^{0}$, formalizing the forward process as
@@
where $F_{p}(\cdot)$ denotes the prior predictor and $\hat{\alpha}_{t}=1-t/T$. The reverse process initializes at $\mathbf{x}_{1:H}^{T}=\mathbf{h}$ and iteratively refines predictions using
@@
where $\alpha_{t}$ decreases with $t$, $\mathcal{S}_{w}(\cdot)$ is a evolution trend extraction function as
@@
The reverse process starts from $\mathbf{x}_{-T+1:0}^{H}=\mathcal{S}_{w}(\mathbf{x}_{-T+1:0})$ and predicts $\mathbf{x}_{2-t:T-t+1}^{t-1}$ by
@@
where $\hat{X}^{0}_{\theta}(\cdot)$ is the devolution network and $\hat{z}_{t}$ is the predicted evolution trend. This formulation inherently embeds the historical information as the generative endpoint, ensuring the reverse trajectory learns a direct mapping from history to future, enhancing temporal dependence learning and facilitating sequence modeling.
@@ 5 Datasets
The datasets used for evaluating diffusion-based TSF models are broadly categorized into two types: unimodal time series datasets and multimodal time series datasets. Unimodal time series datasets consist solely of temporal signals, where all available information is encoded in time-indexed numerical sequences. Multimodal datasets integrate time series with other modalities such as text, images, or metadata, enabling richer contextual modeling. Table 1 illustrates representative examples of the four major types of time series data. Table 2 summarizes the benchmark datasets and evaluation metrics used in each diffusion-based TSF approach.
Table 1: Examples of datasets across different time series types.
@@
Spatio-temporal PEMS04 [Uncaptioned image]
@@
Table 2: Summary of benchmark TSF datasets and evaluation metrics employed in diffusion models.
@@
TimeGrad [35] Exchange, Solar, Electricity, Traffic, Taxi, Wikipedia CRPS
CSDI [60] Solar, Electricity, Traffic, Taxi, Wiki CRPS
SSSD [36] PTB-XL, Electricity, ETTm1 MSE, MAE, RMSE, MRE
DYffusion [63] SST, Navier-Stokes CRPS, MSE, SSR, Time
TimeDiff [64] NorPool, Caiso, Weather, ETTm1, Wind, Traffic, Electricity, ETTh1, Exchange MSE, Time
MG-TSD [52] Solar, Electricity, Traffic, KDD-cup, Taxi, Wikipedia CRPS, NMAE, NRMSE
TMDM [61] Exchange, ILI, ETTm2, Electricity, Traffic, Weather MAE, MSE
mr-Diff [50] NorPool, Caiso, Traffic, Electricity, Weather, Exchange, ETTh1, ETTm1, Wind MAE
LDT [46] Solar, Electricity, Traffic, Taxi, Wikipedia CRPS, MSE
StochDiff [59] Exchange, Weather, Electricity, Solar, ECochG, MMG NRMSE, CRPS
Bim-Diff [65] NorPool, Caiso, Electricity, Weather, Exchange, ETTh1, ETTm1, Wind MAE
CCDM [66] ETTh1, Exchange, Weather, Appliance, Electricity, Traffic MSE, CRPS
FDF [51] ETTh1, ETTh2, ETTm1, ETTm2, Electricity, Exchange, Weather, Wind MSE, MAE
REDI [68] Weather, Exchange, M4-Yearly, M4-Quarterly, M4-Monthly MAE, RMSE
RATD [55] Exchange, Wind, Electricity, Weather MSE, MAE, CRPS
S2DBM [69] ETTh1, ETTh2, ETTm1, ETTm2, ILI, Weather, Exchange MSE, MAE
ARMD [45] Metric, Solar, Energy, ETTh1, ETTh2, ETTm1, ETTm2, Exchange, Stock MSE, MAE
D3U [67] ETTm1, ETTm2, Weather, Solar-Energy, Electricity, Traffic MSE, MAE
LDM4TS [57] ETTh1, ETTh2, ETTm1, ETTm2, Weather, ECL, Traffic MSE, MAE
@@
NsDiff [62] ETTh1, ETTh2, ETTm1, ETTm2, ECL, EXG, ILI, Solar, Traffic CRPS, QICE
FALDA [54] ILI, Exchange, ETTm2, Weather, Electricity, Traffic CRPS
@@ 5.1 Unimodal Time Series Datasets
Unimodal time series datasets refer to those where the input consists of purely temporal numerical data, without involving additional modalities such as text or images. These datasets are further categorized into univariate, multivariate, and spatio-temporal types. Univariate sequences track the evolution of a single variable over time, while multivariate ones involve multiple variables with temporal and cross-variable dependencies. In practice, most datasets are inherently multivariate, and univariate sequences are often constructed by selecting a single variable from them. Spatio-temporal sequences further capture both temporal dynamics and spatial dependencies across locations. We summarize several commonly used unimodal time series datasets below.
@@ 5.1.1 Univariate and Multivariate
Exchange Rates [70]: A daily multivariate dataset consisting of the exchange rates of 8 foreign currencies relative to USD. It captures long-term trends and cross-currency correlations, making it suitable for financial forecasting.
Weather  [^5] https://www.bgc-jena.mpg.de/wetter/: A multivariate dataset collected from 21 weather stations in Germany, containing hourly meteorological variables such as temperature, humidity, and wind speed. It is widely used for evaluating long-range forecasting models.
@@
Traffic [^7] https://pems.dot.ca.gov/: A multivariate dataset from the Caltrans Performance Measurement System (PeMS), containing 10-minute interval occupancy rates from 963 traffic sensors in California. It exhibits strong daily patterns and is often used for modeling temporal dependencies across spatial sensors.
@@
ILI (Influenza-Like Illness) [^9] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html: A univariate weekly dataset of U.S. outpatient visits due to influenza-like illness, published by the CDC. It is widely used for epidemic forecasting and seasonal modeling. ETT (Electricity Transformer Temperature) [71]: A multivariate dataset capturing the hourly oil temperature of electricity transformers and related external features (e.g., load, ambient temperature). It contains multiple variants (ETTh1, ETTh2, ETTm1, ETTm2) differing in frequency and span.
M3 (Makridakis Competition 3) [72]: A benchmark collection of 3,003 univariate time series from the third Makridakis forecasting competition (M3), covering yearly, quarterly, monthly, and other frequencies across domains such as finance, economics, and industry. It remains widely used for evaluating univariate forecasting accuracy and generalization ability.

### v2

@@ 4.2 Diffusion-centric Condition
Diffusion-centric approaches modify the forward process and the reverse process of diffusion models for time series. We categorize these approaches into two groups, namely Enhanced process and Reformulated process. The first group injects time series priors into the forward trajectory and reuses these priors in the reverse step. The second group redesigns the forward dynamics to better match the sequence structure.
Enhanced process Real-world time series present time-varying uncertainty due to external factors, whereas the standard diffusion endpoint is a fixed Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$ shared by all time series. To address this mismatch, this group integrates historical priors into the forward noising trajectory and conditions the reverse step on the same priors. For example, TMDM [95] integrates a Transformer to extract historical features $\hat{\mathbf{y}}_{1:H}$ from $\mathbf{x}_{-L+1:0}$, utilizing this as a conditional prior in both diffusion processes. The modified forward process interpolates between the future time series $\mathbf{x}_{1:H}^{0}$ and the conditional prior $\hat{\mathbf{y}}_{1:H}$ as
@@
In TMDM, the endpoint distribution at diffusion step $T$ is $\mathcal{N}(\hat{\boldsymbol{y}}_{1:H},\mathbf{I})$. The reverse process employs the prior $\hat{\mathbf{y}}_{1:H}$ in the denoising function $\epsilon_{\theta}(\cdot)$ and calculate the $\mathbf{X}_{1:H}^{t-1}$ by
$\begin{split}&\mathbf{X}_{1:H}^{t-1}=\\ &\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{1:H}^{t}-(1-\sqrt{\bar{\alpha}_{t}})\hat{\mathbf{y}}_{1:H}-{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{1:H}^{t},\hat{\mathbf{y}}_{1:H},t)\right)\end{split}$ (39)
@@
where $\gamma_{0}$, $\gamma_{1}$, $\gamma_{2}$ and $\tilde{\boldsymbol{\beta}}_{t}$ are parameters that depend only on $\beta_{t}$. This joint conditioning strategy leverages historical dependencies to refine uncertainty estimation in prediction. Building on TMDM, NsDiff [96] further incorporates a historical-dependent variance to model non-stationary uncertainty. The forward process transitions to an endpoint distribution, which is formulated by
@@
$\begin{split}\mathbf{x}_{1:H}^{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{1:H}^{0}+(1-\sqrt{\bar{\alpha}_{t}})\mathbf{f}(\mathbf{x}_{-L+1:0})\\ +\sqrt{(\bar{\beta}_{t}-\tilde{\beta}_{t})\mathbf{g}(\mathbf{x}_{-L+1:0})+\tilde{\beta}_{t}\sigma_{\mathbf{x}_{1:H}}}\epsilon\end{split}$ (42)
where $\bar{\beta}_{t}$ and $\tilde{\beta}_{t}$ are parameters that depend on $\alpha_{t}$ and $\sigma_{\mathbf{x}_{1:H}}$ is the actual variance. The reverse process conditions denoising on both $\mathbf{f}(\mathbf{x}_{-L+1:0})$ and $\mathbf{g}(\mathbf{x}_{-L+1:0})$ to refine uncertainty estimation. This approach adapts the diffusion trajectory to temporal uncertainty shifts, improving distributional flexibility. Besides modeling uncertainty, another study integrates recurrent fusion of historical priors within the forward diffusion process to enhance the utilization of recent temporal information. REDI [98] implements a recurrent diffusion mechanism where each forward step combines the current noisy future time series value $\mathbf{x}_{i}^{t}$ at time step $i$ with the immediate historical value $\mathbf{x}_{i-1}^{t}$ before noise addition:
@@
This recurrent formulation strengthens focus on proximate history to mitigate distribution drift and ensures strict temporal causality by preventing information leakage from future states during generation. MG-TSD [88] and mr-Diff [68] adopt multi-scale forward guidance. They build scale-specific priors $\{\hat{\mathbf{y}}_{1:H}^{(s)}\}_{s=1}^{S}$ by downsampling or smoothing the history and aligning the result to the horizon length. The forward step at scale $s$ interpolates toward the corresponding prior through the following process:
$\mathbf{x}_{1:H}^{t,(s)}=\sqrt{\alpha_{t}}\,\mathbf{x}_{1:H}^{0,(s)}+(1-\sqrt{\alpha_{t}})\,\hat{\mathbf{y}}_{1:H}^{(s)}+\sqrt{1-\alpha_{t}}\,\epsilon$ (45)
where $\mathbf{x}_{1:H}^{t,(s)}$ denotes the horizon state at diffusion step $t$ under scale $s$. The hierarchy proceeds from coarse to fine. The reverse step at scale $s$ conditions on $\hat{\mathbf{y}}_{1:H}^{(s)}$ and on the upsampled coarse estimate $U^{(s)}(\tilde{\mathbf{x}}^{(s-1)})$ to refine the current resolution, which is formulated by
$\mathbf{x}_{1:H}^{t-1,(s)}=\mu_{\theta}\!\big{(}\mathbf{x}_{1:H}^{t,(s)},\,\hat{\mathbf{y}}_{1:H}^{(s)},\,U^{(s)}(\tilde{\mathbf{x}}^{(s-1)}),\,t\big{)}$ (46)
where $U^{(s)}(\cdot)$ denotes an upsampling operator and $\tilde{\mathbf{x}}^{(s-1)}$ denotes the latest coarse estimate. This coarse-to-fine schedule injects multi-scale priors into the forward trajectory and reuses them in the reverse refinement.
Reformulated process This group replaces the standard noise addition with alternative forward dynamics that reflect temporal structure. Building on the above, this line of studies develops new forward formulations that capture the inherent characteristics of time series. For instance, DYffusion [97] leverages the temporal dynamics in the data, directly coupling it with the diffusion steps. Specifically, this approach designs a stochastic temporal interpolator $I(\cdot)$ for the forward process to generate intermediate states between the initial condition $\mathbf{x}_{0}$ and the target future point $\mathbf{x}_{H}$, formalized as minimizing $\mathbb{E}\left\|I_{\phi}(\mathbf{x}_{0},\mathbf{x}_{H},i)-\mathbf{x}_{i}\right\|^{2}$ where $i$ is an intermediate timestep. The reverse process employs a forecaster $F(\cdot)$ to predict $\mathbf{x}_{H}$ from interpolated states, defined as minimizing $\mathbb{E}\left\|F(\mathbf{x}_{i},i)-\mathbf{x}_{H}\right\|^{2}$, which enhances computational efficiency and enables flexible continuous-time sampling trajectories. S2DBM [99] proposes a Brownian Bridge diffusion process to replace the traditional Gaussian diffusion framework. This approach constructs a stochastic bridge between a historical prior $\mathbf{h}=F_{p}(\mathbf{x}_{-L+1:0})$ and the target future $\mathbf{x}_{1:H}^{0}$, formalizing the forward process as
@@
where $F_{p}(\cdot)$ denotes the prior predictor and $\hat{\alpha}_{t}=1-t/T$. The reverse process initializes at $\mathbf{x}_{1:H}^{T}=\mathbf{h}$ and iteratively refines predictions using the following process:
@@
where $\alpha_{t}$ decreases with $t$, and $\mathcal{S}_{w}(\cdot)$ is a evolution trend extraction function that is formulated as
@@
The reverse process starts from $\mathbf{x}_{-T+1:0}^{H}=\mathcal{S}_{w}(\mathbf{x}_{-T+1:0})$ and predicts $\mathbf{x}_{2-t:T-t+1}^{t-1}$ through the following process:
@@
where $\hat{X}^{0}_{\theta}(\cdot)$ is the devolution network and $\hat{z}_{t}$ is the predicted evolution trend. This formulation inherently embeds historical information as the generative endpoint, ensuring the reverse trajectory learns a direct mapping from the history to the future, enhancing temporal dependence learning and facilitating sequence modeling.
@@ 5 Datasets
The datasets used for evaluating diffusion-based TSF models are broadly categorized into two types: unimodal time series datasets and multimodal time series datasets. Unimodal time series datasets consist solely of temporal signals, where all available information is encoded in time-indexed numerical sequences. Multimodal datasets integrate time series with other modalities such as text, images, or metadata, enabling richer contextual modeling. Table I presents representative examples of unimodal time series data, while Table II presents representative examples of multimodal time series data. Table III summarizes the benchmark datasets and evaluation metrics used in each diffusion-based TSF approach.
TABLE I: Examples of unimodal time series datasets covering univariate, multivariate, and spatio-temporal types. For the Traffic dataset, a univariate sequence is extracted from the multivariate records by selecting Sensor 0. For the ETTh1 dataset, three representative variables are plotted to illustrate its multivariate structure. For the SST dataset, the temporal series is drawn at four spatial grid points, each denoted by latitude and longitude coordinates.
@@
Spatio-temporal SST [Uncaptioned image]
TABLE II: Examples of multimodal TSF datasets. The TTC dataset provides temporally aligned pairs where each timestamp contains both multivariate time series values and their corresponding textual description. The Image-EEG dataset aligns neural time series signals with visual stimuli, where each time series variable is associated with a corresponding image.
Type Datasets Example
@@
Multimodal Image-EEG [Uncaptioned image]
TABLE III: Summary of benchmark TSF datasets and evaluation metrics employed in diffusion models.
@@
CSDI [63] Solar, Electricity, Traffic CRPS
SSSD [64] Electricity, ETTm1 MSE, MAE, RMSE, MRE
DYffusion [97] Navier-Stokes Flow CRPS, MSE
TimeDiff [84] Weather, ETTm1, Traffic, Electricity, ETTh1, Exchange Rates MSE
MG-TSD [88] Solar Energy, Electricity, Traffic CRPS, NMAE, NRMSE
TMDM [95] Exchange Rates, ILI, ETTm2, Electricity, Traffic, Weather CRPS, MAE, MSE
mr-Diff [68] Traffic, Electricity, Weather, Exchange Rates, ETTh1, ETTm1 MAE
LDT [80] Solar Energy, Electricity, Traffic CRPS, MSE
StochDiff [86] Exchange Rates, Weather, Electricity, Solar Energy NRMSE, CRPS
Bim-Diff [85] Electricity, Weather, Exchange, ETTh1, ETTm1 MAE
CCDM [87] ETTh1, Exchange Rates, Weather, Electricity, Traffic MSE, CRPS
FDF [89] ETTh1, ETTh2, ETTm1, ETTm2, Electricity, Exchange Rates, Weather, Wind MSE, MAE
REDI [98] Weather, Exchange Rates MAE, RMSE
RATD [67] Exchange Rates, Electricity, Weather MSE, MAE, CRPS
S2DBM [99] ETTh1, ETTh2, ETTm1, ETTm2, ILI, Weather, Exchange Rates MSE, MAE
ARMD [81] Metric, Solar Energy, ETTh1, ETTh2, ETTm1, ETTm2, Exchange Rates MSE, MAE
D3U [69] ETTm1, ETTm2, Weather, Solar Energy, Electricity, Traffic MSE, MAE
LDM4TS [93] ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, Traffic MSE, MAE
@@
NsDiff [96] ETTh1, ETTh2, ETTm1, ETTm2, Electricity, Exchange Rates, ILI, Solar, Traffic MSE, MAE, CRPS
FALDA [91] ILI, Exchange, ETTm2, Weather, Electricity, Traffic MSE, MAE, CRPS
Diffusion-TS [66] ETTh, Solar Energy MAE, MSE
@@ 5.1 Unimodal Time Series Datasets
Unimodal time series datasets refer to those where the input consists of purely temporal numerical data, without involving additional modalities such as text or images. These datasets are further categorized into univariate [^4] Univariate time series are constructed by decomposing a multivariate series into separate dimensions, each for an individual variable., multivariate, and spatio-temporal types. Univariate sequences track the evolution of a single variable over time, while multivariate ones involve multiple variables with temporal and cross-variable dependencies. In practice, most datasets are inherently multivariate, and univariate sequences are often constructed by selecting a single variable from them. Spatio-temporal sequences further capture both temporal dynamics and spatial dependencies across locations. We summarize several commonly used unimodal time series datasets below.
@@ 5.1.1 Univariate and Multivariate
  • Exchange Rates [^5] https://github.com/datasets/exchange-rates/tree/main[100]: A daily multivariate dataset consisting of the exchange rates of eight foreign currencies relative to US dollars. It captures long-term trends and cross-currency correlations, making it suitable for the TSF task in the financial domain.
  • Weather [^6] https://www.bgc-jena.mpg.de/wetter/: A multivariate dataset collected from 21 weather stations in Germany, containing hourly meteorological variables such as temperature, humidity, and wind speed. It is widely used for evaluating the model performance for long-range TSF.
@@
  • Traffic [^8] https://pems.dot.ca.gov./?dnode=Clearinghouse: A multivariate dataset from the Caltrans Performance Measurement System (PeMS) [^9] https://pems.dot.ca.gov/, containing 10-minute interval occupancy rates from 963 traffic sensors in California. It exhibits strong daily patterns and is often used for training and evaluating approaches to modeling temporal dependencies across spatial sensors.
@@
  • ILI (Influenza-Like Illness) [^11] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html: A multivariate weekly dataset of U.S. outpatient visits due to influenza-like illness, published by the CDC. It is widely used for epidemic forecasting and seasonal modeling.
  • ETT (Electricity Transformer Temperature) [^12] https://github.com/zhouhaoyi/Informer2020[101]: A multivariate dataset capturing the hourly oil temperature of electricity transformers and related external features (e.g., load, ambient temperature). It contains multiple variants (ETTh1, ETTh2, ETTm1, ETTm2) differing in frequency and span.
  • M3 (Makridakis Competition 3) [^13] https://github.com/jordicolomer/m3-competition/blob/master/M3Forecast.xls[102]: A benchmark collection of 3,003 univariate time series from the third Makridakis forecasting competition (M3), covering yearly, quarterly, monthly, and other frequencies across domains such as finance, economics, and industry. It remains widely used for evaluating univariate forecasting accuracy and generalization ability.